OpenStudy (anonymous):

Evaluate the following expression: 2/6+i

OpenStudy (anonymous):

A. \[\frac{ 12 }{ 37 } + \frac{ 2 }{ 37 } i\] B. \[\frac{ 12 }{ 35 } - \frac{ 2 }{ 35 }i\] C. \[\frac{ 12 }{ 35 } + \frac{ 2 }{ 35 }i\] D. \[\frac{ 12 }{ 37 } - \frac{ 2 }{ 37 }i\]

OpenStudy (anonymous):

@dmndlife24

OpenStudy (anonymous):

 solve this